i dont understand how to do equations, please help? t/5-2=6, -2y-5=-9, d/-3+7= 10
Answer by stanbon(75887)

t/5-2=6
Add 2 to both sides to get:
t/5 = 8
Multiply both sides by 5 to get:
t = 40
============================================
-2y-5=-9
Add 5 to both sides:
-2y = -4
Divide by sides by -2:
y = 2
============================================

d/-3 + 7 = 10
Subtract 7 from both sides:
d/-3 = 3
Multiply both sides by -3 to get:
d = -9
============================================
Cheers,
Stan H.
============================